# InterQnet: A Heterogeneous Full-Stack Approach to Co-designing Scalable Quantum Networks

### v1

@@ Abstract
Keywords InterQnet $\cdot$ heterogeneous $\cdot$ scalable $\cdot$ quantum networks $\cdot$ systems approach
1 Introduction
@@
Yet despite the demonstration of entanglement distribution and teleportation of quantum states in laboratory-, campus-, and metropolitan-scale networks [3], scaling quantum networks to longer distances and to multiple users and applications remains a major challenge. The InterQnet project is advancing the state of the art in scalable quantum networks through work aimed at two interrelated overarching goals: (1) practical realization of a full-stack heterogeneous quantum network (InterQnet-Achieve) and (2) systems study of scalable quantum network architectures (InterQnet-Scale). By accomplishing these goals, InterQnet will provide a fundamental building block and foundational architectural choices for national-scale quantum networks.
Figure 1: Our systems approach to scalable quantum networks. The approach integrates devices, error correction methods, protocols, architectures, and simulation/experimentation into two co-design/integration cycles.
@@
The remainder of the paper is organized as follows. Section 2 reviews the most recent experimental demonstrations of quantum networks and proposals of architectures for quantum networks, highlighting the main challenges for scalability. Section 3 describes our approach to address scalability challenges. As illustrated in Fig. 1, InterQnet involves three thrusts: Heterogeneous Devices (Section 4), Error Management (Section 5), and Network Architecture and Protocols (Section 6). These sections, along with Section 7, provide details on our progress toward the InterQnet goals. Section 8 concludes the manuscript with an outlook for future achievements.
2 Background
@@
Researchers have demonstrated early prototypes of campus- and metropolitan-scale quantum communications [5, 6, 7, 8]. However, unavoidable signal losses make quantum repeaters essential for long-range ($\gtrsim$ 100-kilometer) communication, with repeater-assisted entanglement distribution envisioned through quantum memories [9]. Meanwhile, on the theory front, several quantum network architectures and protocol stacks have been proposed for a future quantum internet [10]. For instance, Van Meter et al. [11] proposed a layered stack approach and later developed this into a recursive architecture [12]. Wehner’s group at TU Delft has also proposed layered models for quantum networks [13, 14] that roughly map the classical TCP/IP stack. While Van Meter’s and Wehner’s work focused on bipartite entangled states, Pirker and Dür [15] extended this consideration to multipartite entangled states. Despite these advances, quantum networking still has a long way to go before it becomes a mature technology. A significant gap persists between the theory and practice of quantum networking, and one could argue that this gap may even be widening. The following subsections provide details on efforts that attempt to close this gap by building a quantum networking testbed and defining novel quantum networking architectures.
2.1 Quantum Networking Testbeds
@@
Recent efforts to build and characterize deployed-fiber quantum network testbeds are making progress in bridging the gap between theory and practical implementation. Pompili et al. demonstrated a phase-stabilized quantum network architecture based on three entangled nodes [16] and achieved entanglement delivery [7] using the protocol stack proposed in [13]. The BARQNET testbed [17] has provided valuable insights into characterizing and compensating for channel impairments including polarization and phase noise over a 50 km fiber loop, highlighting effective integration strategies for quantum memory compatibility. The QUANT-NET testbed [18, 19] building a multinode quantum network connecting Lawrence Berkeley National Laboratory and UC Berkeley over deployed fiber, with the goal of implementing entanglement swapping and teleportation of controlled-NOT gates between distant trapped-ion processors. In the interim, QUANT-NET has deployed essential infrastructure components such as Hong–Ou–Mandel interferometry and polarization drift compensation, controlled by a two-level control framework [20].
2.2 Quantum Network Architectures
@@
Early attempts to define an architecture and protocol stack for a quantum internet have tried to replicate the success of the classical Internet; thus, their proponents defined stacks [21, 13, 6] that map directly to the TCP/IP [22] stack. Delle Donne et al. [23] presented QNodeOS, a platform-independent operating system for end nodes of a quantum network, which enables multitasking and high-level application execution on heterogeneous quantum hardware. Besides these architectures, researchers have proposed software-defined networking-based control for quantum networks [24, 25, 26].
@@
2.3 Open Scalability Challenges
Figure 2: Hourglass-shaped quantum networking stack. EPS is an entangled photon source.
@@ 2.3 Open Scalability Challenges
The transition from small-scale, isolated experiments to national-scale, heterogeneous quantum networks is hindered by limitations in scalability, control, and architectural understanding. Most current efforts are restricted to homogeneous platforms and linear topologies, with little validation against real-world conditions.
Scalability in quantum networks is multidimensional. Distance remains limited by photon loss and memory decoherence. Equally important is interoperability, since practical networks must link superconducting qubits, atomic ensembles, and solid-state memories operating at different wavelengths and bandwidths. Entanglement distribution rates are constrained by probabilistic sources, multipair emissions, and imperfect detectors. Supporting multiple simultaneous users introduces new demands for orchestration and fair allocation of entangled pairs. Resilience is also critical, as accumulated noise threatens fidelity without effective error management.
Critical issues remain. Today’s networks lack well-defined abstractions for component classes, leaving it unclear how to build robust control layers across diverse technologies. Error correction and mitigation must be deployed at each node, customized to the service being provided. Architecture-level design choices are rarely guided by realistic performance and resource constraints. Interfaces that allow users and applications to access quantum network services remain poorly defined. It is also unclear how traffic engineering and quality of service can be realized when photonic states cannot be directly processed in the network.
@@
A useful lesson from the classical Internet is not the TCP/IP stack itself but its hourglass shape. A single narrow waist connects many protocols above and below. In quantum networks, entanglement protocols can play this role, accommodating heterogeneous physical platforms underneath and diverse applications on top (see Fig. 2). Achieving such an hourglass-shaped stack will require clear abstractions for components, scalable error management, and a transversal layer for management, monitoring, and validation.
@@
These open challenges motivate the dual goals of InterQnet-Achieve and InterQnet-Scale described in Section 3. Sections 4 through 7 then describe our approach: heterogeneous devices, error management, network architecture, and integration.
3 The InterQnet Approach
@@
Quantum communication networks must scale along multiple dimensions, including heterogeneity of qubit platforms, distance between quantum information source and destination, number of quantum nodes in the network, rate of teleportation or entanglement distribution, and number of users. InterQnet takes a full-stack, co-design approach to enable scalable quantum networks through two overarching goals:
@@
  • InterQnet-Achieve: Practical realization of a 3-node heterogeneous quantum network that integrates ytterbium atoms, erbium ions, and superconducting qubits, connected over the Argonne quantum network (ARQNET) [29] fiber infrastructure with quantum frequency conversion and microwave-optical transduction.
@@
  • InterQnet-Scale: A systems study focused on architectural trade-offs critical to scalability, including centralized vs. distributed control, in-band vs. out-of-band signaling, and resource management across heterogeneous platforms.
@@
4 Heterogeneous Devices
A quantum repeater is a key element in a scalable quantum network. Component devices such as quantum memories and quantum frequency converters are essential to the realization of a quantum repeater. We leverage three different types of state-of-the-art qubit systems that are being developed at Argonne as quantum nodes to demonstrate a quantum network repeater (Fig. 3): trapped ytterbium (Yb) atom arrays, erbium ions (Er3+) in solids, and superconducting quantum devices. This heterogeneous architecture addresses a key scalability requirement by allowing for remote quantum entanglement and communication between different kinds of quantum information carriers in both the optical and microwave regimes.
Each of the three qubit systems can generate either photon-photon or photon-spin entanglement to serve as a quantum node for entanglement swapping and entanglement-based quantum communications. We will use the Yb atom array system as the key quantum repeater station; this will provide both long-lived quantum memories based on the atom spin states and the necessary two-qubit gates between local atoms for implementing the error mitigation schemes of Thrust 2 (see Section 5). Below, we provide detailed descriptions of the three distinct qubit platforms and clarify the role of each component device.
4.1 Ytterbium atoms
@@ 4.2 Erbium Ions
An optical image of the Er3+ device is shown in Fig. 3(b). The device is a thin film of Er-doped titanium dioxide (TiO2) with low background spins, directly grown on top of an silicon-on-insulator (SOI) substrate. Fabricating 1D nanophotonic crystal cavities to couple with the Er3+ results in an optical photon emission Purcell enhancement factor exceeding 200, as well as a high single-photon collection efficiency ($\sim$ 40%) via fiber-to-chip coupling [42]. Similar devices of Er3+ in TiO2 thin films grown post-fabrication on SiN waveguides boast a narrow ensemble homogeneous linewidth of 5 kHz and spectral diffusion of 27 kHz over 4 ms, well within the 1 MHz microwave-optical (M-O) transducer bandwidth [43].
@@
4.3 Superconducting quantum devices
A scalable quantum communication network should be able to not only transmit quantum information carried by optical photons but also interface with solid-state qubits working at the microwave frequencies [10, 46]. At the microwave quantum node, we will leverage GHz qubit devices based on superconducting transmons and/or single electron-on-solid neon (eNe) qubits [47] to create and receive microwave quantum communication signals and integrated quantum transducers to generate M-O photon entanglement [48, 49], hence achieving remote entanglement between a microwave qubit and a distant Yb atom or Er ion.
@@ 4.3 Superconducting quantum devices
4.4 Optical Quantum Frequency Converters
@@
We have designed QFC schemes in periodically poled lithium niobate waveguides with a strong $\chi^{(2)}$ nonlinearity (see Fig. 3(d)). We focus on bulk-optic designs to keep coupling losses low and will use cavity enhancement to match the interaction bandwidth to those of the incident photons. By using pump wavelengths in the telecom bands, we will be able to convert any pair of wavelengths specified above to the common wavelength of 745.7 nm through sum frequency generation. Besides bandwidth, conversion efficiency, and added noise, another important co-design parameter is the temporal shape of the single photons, which needs to be matched between different quantum emitters to remove photon distinguishability. We have previously shown that such wavefunction engineering can be done by controlling the temporal shape of the pump beam [57, 58] during quantum frequency conversion. The capability of temporal mode shaping will play an important role in our heterogeneous network, allowing us to maximize visibility of two-photon interference and hence optimize Bell state measurements.
4.5 Summary of Achievements and Open Challenges

### v2

@@ Abstract
I Introduction
@@
Yet despite the demonstration of entanglement distribution and teleportation of quantum states in laboratory-, campus-, and metropolitan-scale networks [76], scaling quantum networks to longer distances and to multiple users and applications remains a major challenge. The InterQnet project is advancing the state of the art in scalable quantum networks through work aimed at two interrelated overarching goals: (1) practical realization of a full-stack heterogeneous quantum network (InterQnet-Achieve) and (2) systems study of scalable quantum network architectures (InterQnet-Scale). By accomplishing these goals, InterQnet will provide fundamental building blocks and foundational architectural choices for national-scale quantum networks.
@@
Figure 1: Organization of the InterQnet project, highlighting our systems approach to scalable quantum networks. The approach integrates devices, error correction methods, protocols, architectures, and simulation/experimentation into two co-design/integration cycles.
@@
The remainder of the paper is organized as follows. Section II reviews the most recent experimental demonstrations of quantum networks and proposals of architectures for quantum networks, highlighting the main challenges for scalability. Section III describes our approach to address scalability challenges. As illustrated in Fig. 1, InterQnet involves three thrusts: Heterogeneous Devices (Section IV), Error Management (Section V), and Network Architecture and Protocols (Section VI). These sections, along with Section VII, provide details on our progress toward the InterQnet goals. Section VIII concludes the manuscript with an outlook for future achievements.
II Background
@@
Researchers have demonstrated early prototypes of campus- and metropolitan-scale quantum communications [120, 2, 98, 119, 116, 60, 75, 27]. However, unavoidable signal losses make quantum repeaters essential for long-range ($\gtrsim$ 100-kilometer) communication, with repeater-assisted entanglement distribution envisioned through quantum memories [66]. Meanwhile, on the theory front, several quantum network architectures and protocol stacks have been proposed for a future quantum internet [59]. For instance, Van Meter et al. [122] proposed a layered stack approach and later developed this into a recursive architecture [123]. Wehner’s group at TU Delft has also proposed layered models for quantum networks [30, 64] that roughly map the classical TCP/IP stack. While Van Meter’s and Wehner’s work focused on the distribution of bipartite entangled states on demand, Pirker and Dür [96] envisioned an architecture in which multipartite entangled states are distributed in advance and consumed in response to user requests. However, early architecture proposals for a quantum internet impose stringent performance requirements on the underlying hardware devices. Thus, a significant gap persists between the theory and practice of quantum networking. The following subsections provide details on efforts that attempt to close this gap by building a quantum networking testbed, defining novel quantum networking architectures, and evaluating architectural choices via simulations.
II-A Quantum Networking Testbeds
@@
Recent efforts to build and characterize deployed-fiber quantum network testbeds are making progress in bridging the gap between theory and practical implementation. Pompili et al. demonstrated a phase-stabilized quantum network architecture based on three entangled nodes [99] and achieved entanglement delivery [98] using the protocol stack proposed in [30]. Similarly, Stolk et al. [116] demonstrated heralded entanglement generation of predefined states between two diamond spin nodes in Delft and The Hague. The two nodes were connected to a midpoint station over a total distance of 25 km of deployed fiber, and photons were converted to the telecom L-band. Knaut et al. [60] demonstrated nuclear spin entanglement distribution between two SiV nodes connected through 35 km of deployed fiber routed across the greater Boston metropolitan area, with quantum frequency conversion to the telecom O-band. Similarly, Liu et al. [75] demonstrated the creation of entanglement between two rubidium quantum memories separated by 12.5 km of deployed fiber in the city of Hefei. Craddock et al. [27] used the GothamQ testbed in New York to demonstrate automated distribution of polarization-entangled photons over 34 km of underground deployed fiber. The authors reported continuous distribution of entangled pairs for 15 days, with a network uptime of 99.84%. The BARQNET testbed [12] has provided valuable insights into characterizing and compensating for channel impairments including polarization and phase noise over a 50 km fiber loop, highlighting effective integration strategies for quantum memory compatibility. The QUANT-NET testbed [83, 110] is building a multinode quantum network connecting Lawrence Berkeley National Laboratory and UC Berkeley over deployed fiber, to implement entanglement swapping and teleportation of controlled-NOT gates between distant trapped-ion processors. In the interim, QUANT-NET has deployed essential infrastructure components such as Hong–Ou–Mandel interferometry and polarization drift compensation, controlled by a two-level control framework [130].
II-B Quantum Network Architectures
@@
Early attempts to define an architecture and protocol stack for a quantum internet have tried to replicate the success of the classical Internet; thus, their proponents defined stacks [95, 30, 2] that map directly to the TCP/IP [40] stack. As mentioned before, these architectures impose stringent requirements on hardware devices. These requirements include asynchronous operations enabled by quantum memories with long coherence times and gate operations and measurements with high fidelity [123] so protocols can run in a distributed fashion. Taking a step toward TCP/IP-inspired architectures, Delle Donne et al. [34] presented QNodeOS, a platform-independent operating system for end nodes of a quantum network. QNodeOS enables multitasking and high-level application execution on heterogeneous quantum hardware. However, current near-term quantum networks—of campus and metropolitan scale—will require centralized control and architectures that follow software-defined networking principles [80, 23, 10]. Moreover, the current consensus is that these networks will serve entanglement requests as they arrive at the central controller. The QUANT-NET [130] and Quantum Internet Alliance [9] groups have demonstrated implementations of this approach, and we adopt a centralized control architecture for our InterQnet-Achieve demonstration.
While on-demand entanglement distribution may become the architecture for long-distance quantum networks, architectures that continuously generate link-level entanglement deserve attention. In the context of distributed quantum computing, Talsma et al. [118] analyzed a protocol for continuous entanglement distribution among nodes arranged in regular patterns (e.g., linear, honeycomb, square grid, and triangle lattice topologies). The study found that the number of swap operations required to increase the virtual neighborhood of a node (i.e., the number of entangled neighbors of a given node) increases as the coherence time of quantum memories decreases. Gu et al. [46] proposed a framework called ESDI for long-distance entanglement scheduling and distribution in a general quantum network. By leveraging discrete-time simulations, the authors showed that ESDI can significantly reduce the average completion time of long-distance entanglement requests. Vardoyan et al. [124] studied the capacity region of a switch that can generate bipartite or tripartite entanglement between end users connected to the switch in a star topology. The study showed that a set of randomized switching policies can outperform a time-division multiplexing policy, even when quantum memory decoherence is considered.
@@
II-C Simulations of Quantum Networks
Simulators have been used to study architectural design choices of quantum networks, with repeater chains being a popular subject of study in the theory community. For instance, Avis et al. [6] studied the hardware requirements of a quantum repeater node deployed on a hypothetical metropolitan-scale fiber network. The network connects the cities of Delft and Eindhoven (separated by 226.5 km), with the repeater located in the city of Utrecht. Besides realistic parameters of the fiber network, the authors considered the parameters for repeaters based on color centers in diamonds and trapped ion devices. Their results confirm the discrepancies between abstract/simplified and more realistic models. Moreover, the simulation results provide minimal hardware requirements to build quantum repeaters based on the chosen hardware platforms. Similarly, Jain et al. [57] created accurate models of trapped ion quantum repeaters and simulated the performance of repeater chains using current and future parameters of the devices. The authors considered a hypothetical deployment over the ESnet network—a research and education network in the United States—and concluded that current trapped-ion platforms are unable to yield any usable entanglement fidelities and rates. Thus, device improvements, purification schemes, and quantum error correction mechanisms are needed for these platforms to become practical.
Other architectural design choices have been evaluated via simulations. For example, Chan et al. [20] studied chains of all-photonic quantum repeaters. Soon et al. [115] investigated quantum networks composed of heterogeneous links (i.e., some links used the meet-in-the-middle protocol while others used the midpoint source protocol). Their simulations showed that while performance is dependent on link configuration, no significant degradation was observed when compared with homogeneous networks. Koyama et al. [62] investigated the switched interconnects required to build modular quantum computers. The authors proposed designs for reducing loss and crosstalk while raising entanglement rates and fidelity. Van Dam [121] et al. investigated the hardware requirements for executing verifiable blind quantum computation (BQC) using a trapped ion server and a distant measurement-only client, separated by a 50 km link.
II-D Open Scalability Challenges
Figure 2: An entanglement-centric, hourglass-like view of quantum networking. This abstraction is provided for illustrative purposes and will not be evaluated in this work.
@@ II-D Open Scalability Challenges
The transition from small-scale, isolated experiments to national-scale, heterogeneous quantum networks is hindered by limitations in scalability, control, and architectural understanding.
C1. Integration of heterogeneous technologies: Different qubit platforms excel at specific tasks, while having deficiencies in others. Thus, integrating complementary qubit platforms into heterogeneous systems is a promising path toward scalability. Furthermore, the ability to convert between different degrees of freedom in encoded photons [77] and leverage diverse physical media (e.g., fiber optics and free space) [111] will greatly improve heterogeneous quantum networks. Other groups besides InterQnet are also focusing on integrating heterogeneous nodes into their network testbeds. For instance, the QUANT-NET team is planning to demonstrate quantum state teleportation from a G-center node to a trapped-ion node [110]. Similarly, the Quantum Internet Alliance is planning to interconnect trapped-ion end-nodes through a chain of rare-earth quantum repeaters [28]. As seen in the literature review, quantum frequency conversion (QFC) is an important component for interconnecting matter-based qubit platforms over telecom fiber [116, 60]. QFC is already challenging, but the difficulty increases when interconnecting heterogeneous platforms because one now must also consider photon waveform matching [28]. In summary, interconnecting heterogeneous platforms over quantum networks with rates and fidelity sufficiently high for practical applications remains an open challenge.
C2. Control Frameworks for Quantum Networks: While long-term visions of a quantum internet propose architectures that resemble the TCP/IP stack and leverage distributed control [123, 63], recent testbed demonstrations focus on centralized control architectures [130, 9]. Based on their experience implementing and evaluating a modular control architecture for quantum networks, Beauchamp et al. [8] presented a collection of open questions and challenges for developing future quantum network control architectures and proposed potential solutions. Worth highlighting are the need for good admission control for user requests, computationally fast scheduling methods, and the need to know when to establish service agreements between the network and users. Hayek et al. [50] presented a review of software for designing and operating quantum networks, with a special focus on the network functions of the control plane. The authors identified areas that need further attention, such as topology management functions that can handle dynamic reconfiguration, standard interfaces for interoperability, and a management plane for resilience of the network. In summary, advancing the control frameworks of quantum networks from automation of experiments to platforms that actually provide services to users remains an open challenge.
@@
C3. Refined Architectural Understanding: Li et al. [74] presented a comprehensive survey on protocols for a quantum internet, from the perspective of layered models. The authors identified several challenges, such as increasing the realism of models for physical devices, establishing clear boundaries between layers—especially between physical and link, as well as between network and transport, and integrating quantum networks with existing infrastructure. The survey further emphasized the lack of consensus among layered model proposals. We argue that a useful lesson from the classical Internet is not the TCP/IP stack itself but its hourglass shape. A single narrow waist connects many protocols above and below. In quantum networks, entanglement protocols can play this role, accommodating heterogeneous physical platforms underneath and diverse applications on top (see Fig. 2). Achieving such an hourglass-shaped stack will require clear abstractions for components; scalable error management; and a transversal layer for management, monitoring, and validation.
@@
These open challenges motivate the dual goals of InterQnet-Achieve and InterQnet-Scale described in Section III. Sections IV through VII then describe our approach: heterogeneous devices, error management, network architecture, and integration.
III The InterQnet Approach
@@
Quantum communication networks must scale along multiple dimensions, including heterogeneity of qubit platforms, distance between quantum information source and destination, number of quantum nodes in the network, rate of teleportation or entanglement distribution, and number of users. As elaborated in Section II-D, open scalability challenges include the integration of heterogeneous technologies (C1), the development of control frameworks (C2), and a refined understanding of architectures (C3) for quantum networks. InterQnet takes a full-stack, co-design approach to enable scalable quantum networks through two overarching goals:
@@
  • InterQnet-Achieve: Practical realization of a 3-node heterogeneous quantum network that integrates ytterbium atoms, erbium ions, and superconducting qubits, connected over the Argonne quantum network (ARQNET) [56] fiber infrastructure with quantum frequency conversion and microwave-optical transduction. InterQnet-Achieve aims to tackle challenge C1 by conducting heterogeneous device development and integration, with experimental demonstrations of entanglement distribution. Moreover, InterQnet-Achieve will address challenge C2 by adopting a centralized control architecture following SDN design principles.
@@
  • InterQnet-Scale: A systems study focused on architectural trade-offs critical to scalability, including centralized vs. distributed control, in-band vs. out-of-band signaling, and resource management across heterogeneous platforms. InterQnet-Scale aims to close the knowledge gap identified in C3 by conducting modeling and simulation studies considering models informed by experimental data and large-scale network topologies.
@@
IV Heterogeneous Devices
A quantum repeater is a key element in a scalable quantum network. Component devices such as quantum memories and quantum frequency converters are essential to the realization of a quantum repeater. We leverage three types of state-of-the-art qubit systems that are being developed at Argonne as quantum nodes to demonstrate a quantum network repeater (Fig. 3) for InterQnet-Achieve: trapped ytterbium atom arrays, erbium ions (Er3+) in solids, and superconducting quantum devices. This heterogeneous architecture addresses a key scalability requirement (see C1 in Section II-D) by allowing remote quantum entanglement and communication between different kinds of quantum information carriers in both the optical and microwave regimes.
Each of the three qubit systems can generate either photon-photon or photon-spin entanglement to serve as a quantum node for entanglement swapping and entanglement-based quantum communications. Link-level entanglement between heterogeneous devices will be generated by using the meet-in-the-middle protocol with QFC at a Bell state measurement (BSM) node. We will use the Yb atom array system as the key quantum repeater station; this will provide both long-lived quantum memories based on the atom spin states and the necessary two-qubit gates between local atoms for implementing the error mitigation schemes of Thrust 2 (see Section V). Below, we provide detailed descriptions of the three distinct qubit platforms and clarify the role of each component device.
IV-A Ytterbium atoms
@@ IV-B Erbium Ions
An optical image of the Er3+ device is shown in Fig. 3(b). The device is a thin film of Er-doped titanium dioxide (TiO2) with low background spins, directly grown on top of a silicon-on-insulator (SOI) substrate. Fabricating 1D nanophotonic crystal cavities to couple with the Er3+ results in an optical photon emission Purcell enhancement factor exceeding 200, as well as a high single-photon collection efficiency ($\sim$ 40%) via fiber-to-chip coupling [36]. Similar devices of Er3+ in TiO2 thin films grown post-fabrication on SiN waveguides boast a narrow ensemble homogeneous linewidth of 5 kHz and spectral diffusion of 27 kHz over 4 ms, well within the 1 MHz microwave-optical (M-O) transducer bandwidth [47].
@@
IV-C Superconducting quantum devices
A scalable quantum communication network should be able to not only transmit quantum information carried by optical photons but also interface with solid-state qubits working at the microwave frequencies [59, 32] for computation tasks. At the microwave quantum node, we will leverage GHz qubit devices based on superconducting transmons and/or single electron-on-solid neon (eNe) qubits [134] to create and receive microwave quantum communication signals and integrated quantum transducers to generate M-O photon entanglement [106, 81], hence achieving remote entanglement between a microwave qubit and a distant Yb atom or Er ion.
@@ IV-C Superconducting quantum devices
IV-D Optical Quantum Frequency Converters
@@
Key roadblocks to demonstrating high-efficiency, low-noise QFCs are noise processes, such as Raman scattering and weakly phase-matched parametric down-conversion, which arise in nonlinear media due to the strong pump laser field that is required to achieve high-efficiency frequency conversion of photonic quantum states via sum or difference frequency generation. Recent advances in optical to telecom conversion have demonstrated that such noise can be dramatically reduced by limiting the output bandwidth of the QFC to a few megahertz since the noise processes are inherently broadband; however, overall device efficiency suffers because of a reliance on low transmission filtering elements [126].
Our approach utilizes commercial ultra-narrow ($\sim 100$ GHz), high transmission ($95\%$), optical band-pass filters combined with modest finesse ($\mathcal{F}=200$) Fabry Perot optical cavities of multiple lengths to transmit only a single mode of the longest cavity. The bandwidth of the transmitted light can be tuned to match a quantum networking device by properly selecting the length and finesse of the longest cavity. For example, to convert 1389 nm light emitted from Yb atoms, a 10 cm cavity with $\mathcal{F}=200$ is used to transmit a bandwidth of $\sim 7$ MHz, which is broader than the bandwidth of emitted light ($0.5$ MHz). To maintain high transmission through the cavities, we utilize superpolished mirror substrates and low-loss optical coatings.
@@
Besides bandwidth, conversion efficiency, and added noise, another important co-design parameter is the temporal shape of the single photons, which needs to be matched between different quantum emitters to remove photon distinguishability. We have previously shown that such wavefunction engineering can be done by controlling the temporal shape of the pump beam [53, 79] during quantum frequency conversion. The capability of temporal mode shaping will play an important role in our heterogeneous network, allowing us to maximize visibility of two-photon interference and hence optimize Bell state measurements.
TABLE I: Summary of development status for heterogeneous devices.
Device Application State-of-the-Art InterQnet-Achieve Goal Status Challenges
Ytterbium Atom Tweezer Arrays Entanglement distribution; Repeater; Error correction <1 Hz spin-entangled photon emission rate, Fidelity >0.95 [71] 1 Hz heralded entanglement rate, Fidelity >0.99 Implementing spin-photon entanglement on ARQNET Current system limited by photon collection efficiency. Coupling to optical cavities dramatically improves collection [54].
Erbium Ions in TiO2 thin films Memory Spin-photon entanglement rate of $\sim$1.5 Hz over >15 km of optical fiber in hybrid single-ion devices 1 Hz heralded entanglement rate Single-ion devices fabricated and tested in laboratory Oxygen vacancies incorporated during film growth and induced during fabrication limit optical coherence
Superconducting and Single Electron Qubits Computing; Sensing; Memory Processors integrating >>100 transmon qubits [45, 16], Coherence time $\sim$100 $\mu{s}$ [45, 135] Superconducting and/or electron qubits with cavity frequency matched to M-O transducers in quantum network Device design and initial fabrication completed Frequency matching between qubit readout cavity and the M-O transducer
Microwave-optical (M-O) Transducer M-O photon frequency conversion; Entanglement generation Efficiency $\sim$10% with added noise $\sim$0.2 [106] (Higher efficiency of $\sim$48% was achieved with higher noise and lower bandwidth [51]) High efficiency >50% with low added noise $\lesssim 0.1$ for quantum operations Initial device fabrication completed. Device being characterized and will be further optimized High optical pump power induces heating and noise at low temperature. Multiple frequency/wavelength matching between the transducer and other qubit systems
Optical Quantum Frequency Converters O-O photon frequency conversion Total device efficiency 25%, negligible dark count noise [126] Total device efficiency >50%, negligible dark count noise Design completed, working toward lab demonstration High-efficiency conversion creates noise that must be filtered or eliminated
IV-E Summary of Achievements and Open Challenges